OpenStudy (anonymous):

What are the possible steps involved in solving the equation shown? Check all that apply. 3.5 + 1.2(6.3 – 7x) = 9.38 Add 3.5 and 1.2. Distribute 1.2 to 6.3 and –7x . Combine 6.3 and –7x . Combine 3.5 and 7.56. Subtract 11.06 from both sides. Divide both sides by 9.38.

OpenStudy (anonymous):

Distribute 1.2 to 6.3 and –7x Combine 3.5 and 7.56. Subtract 11.06 from both sides

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

can u hlp me with one more

OpenStudy (anonymous):

Ya I'll try

OpenStudy (anonymous):

And you're welcome!

OpenStudy (anonymous):

A number, n, is added to 15 less than 3 times itself. The result is 101. Which equation can be used to find the value of n? 3n – 15 + n = 101 3n + 15 + n = 101 3n – 15 – n = 101 3n + 15 – n = 101

OpenStudy (anonymous):

C.

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

Because 3 times less would be 3n and less than that would mean the negative sign to both the 15 and n.

OpenStudy (anonymous):

And you're welcome! :)